OpenStudy (anonymous):

a landscape artist is desning a park that will be 12 acres.there will be tree, plants, and benches for people to relax on . he wants to plant sections that are 2/5 acres each. explain in words how you can find out the number of 2/5 sections there will be? how many 2/5 acres will there be? so far the land scape artist knows that he wants to plant 3/4 of one of the 2/5 acre sections . what fractions of an acre will be planted in cherry trees

OpenStudy (anonymous):

12 acres / (2/5 acres) = 60/2 = 30 sections. (Divide by fraction). (2/5 acre) * (3/4 percent for cherry tree) = 3/10 of an acre per section for cherry trees Total acreage for cherry trees= 3/10 * 30 = 9 acres. Didn't check my work.

OpenStudy (anonymous):

thank you so much. i tried before but was so wrong. thank again.